# Warm Rotation, and Principled Training for Geometric and Neuromorphic AI

### 18 Mar 2026 v1

@@ Abstract
Prevailing AI training infrastructure assumes reverse-mode automatic differentiation over IEEE-754 arithmetic. The memory overhead of training relative to inference, optimizer complexity, and structural degradation of geometric properties through training are consequences of this arithmetic substrate. This paper develops an alternative training architecture grounded in three prior results: the Dimensional Type System and Deterministic Memory Management framework [6], which establishes stack-eligible gradient allocation and exact quire accumulation as design-time verifiable properties; the Program Hypergraph [8], which establishes grade preservation through geometric algebra computations as a type-level invariant; and the b-posit 2026 standard [10], which makes posit arithmetic tractable across hardware targets conventionally considered inference-only. Their composition enables depth-independent training memory bounded to approximately twice the inference footprint, grade-preserving weight updates, and exact gradient accumulation, applicable uniformly to loss-function-optimized and spike-timing-dependent neuromorphic models. We introduce Bayesian distillation, a mechanism by which the latent prior structure of a general-purpose model is extracted through the ADM training regime, resolving the data-scarcity bootstrapping problem for domain-specific training. For deployment, we introduce warm rotation, an operational pattern in which an updated model transitions into an active inference pathway without service interruption, with structural correctness formalized through PHG certificates and signed version records. The result is a class of domain-specific AI systems that are smaller and more precise than general-purpose models, continuously adaptive, verifiably correct with respect to the physical structure of their domains, and initializable from existing models.
@@ 1.2 The Three Enabling Results
The second is the Program Hypergraph established in [8]. That work demonstrated that grade in Clifford algebra is a type-level invariant under the DTS framework’s dimensional group structure, that grade inference determines the non-zero entries of the geometric product Cayley table at design time, before any arithmetic occurs, and that PHG saturation semantics correctly formalize multi-way geometric constraints that binary edge graphs cannot represent without information loss. The critical extension for training: the dual-number representation used in forward-mode autodiff augments each primal multivector with a tangent component of the same grade. PHG grade inference applies identically to tangent computations. A grade-2 bivector weight has a grade-2 gradient by the chain rule closure of grade-preserving operations, enforced as a design-time constraint enforced by the type system. Grade corruption through training is a type violation, surfacing in the Lattice language server as the code is written. Numerical drift in IEEE-754 would cause the same corruption to accumulate silently; the type system surfaces it immediately.
The third is the b-posit 2026 standard [10]. The bounded regime field (limited to $\leq 6$ bits) reduces posit decoder hardware cost to 79% less power, 71% less area, and 60% less latency relative to standard posit decoders, achieving cost competitive with IEEE-754 float32 decoders on NPU-class hardware. This is the threshold at which posit arithmetic becomes tractable on edge and embedded hardware targets. Below this threshold, posit arithmetic required data center hardware to justify its precision advantages. At this threshold, posit arithmetic with quire accumulation is a realistic substrate for AI training on devices that also run inference.
@@ 3.4 Bayesian Distillation from General-Purpose Models
Recent empirical work [24] establishes the key precondition: latent Bayesian structure in general-purpose LLMs is accessible through targeted fine-tuning and generalizes across domains. This result, developed independently of the ADM framework, confirms that the prior knowledge absorbed during large-scale pretraining is not merely a surface statistical regularity but a structural property that can be elicited and directed.
@@ 5.1 Grade Preservation Through Training
Proof sketch. The dual-number extension of a grade-$k$ primal is a grade-$k$ tangent. PHG grade inference applies identically to the tangent computation as to the primal. The directional derivative $\langle\nabla_{W}\mathcal{L},v\rangle$ is a scalar (grade-0) result, computed as the quire accumulation of products between the grade-$k$ weight tangent components and the random projection $v$. The gradient $\nabla_{W}\mathcal{L}$ is the reconstruction of the weight-space vector from the scalar projection estimates; under the PHG type system, this reconstruction is constrained to the grade-$k$ subspace. Any component at grade $j\neq k$ is a type violation caught at design time. $\square$
@@ 10. Conclusion
Both loss-function gradient descent over Clifford algebra neural networks and spike-timing-dependent plasticity over neuromorphic spiking networks are instances of this architecture. The warm rotation model and versioning infrastructure defined in Section 4 is designed to apply uniformly to both. A system of interlocking domain models within a managed actor framework brings together the precision of expert model traversal and the compositional reach of agentic coordination, while asserting verifiable results from each model and integrity of communication at every shared inter-actor boundary. We assert that this combination, structurally enforced specialization with formally verified message boundaries, is the property neither pure MoE nor current agentic frameworks can provide simultaneously.
General-purpose language models participate in this system in two distinct roles, neither of which is terminal. As Bayesian distillation sources, they seed domain prior initialization with the breadth of their pretraining, providing a tractable starting point where domain-specific data is scarce. As active participants in the running system, they serve interim roles during periods when specialized domain actors are still being developed, trained, or refined, contributing general reasoning capability within the dimensional boundaries the actor framework enforces at the message fabric level. As domain expertise matures within the closed system, general-purpose participants yield progressively more of their roles to structurally verified domain actors, but they remain available as a reasoning resource for problem spaces that do not yet warrant or support a dedicated domain model.
The phrase “ambient AI you can trust” describes a structural property: trust established at design time, through elaboration of the program’s structural properties, grounded in cryptographically anchored version records rather than in benchmark performance against a curated test set.
@@ Acknowledgments
The author thanks John L. Gustafson for guidance on posit arithmetic, the b-posit 2026 standardization pathway, and for Every Bit Counts: Posit Computing [9], whose Chapter 13 provides the definitive parameterization reference for application-specific posit systems. The treatment of representation selection for training arithmetic in Sections 3 and 4 reflects his technical influence directly. The author thanks Barak A. Pearlmutter for his encouragement of this research program and for his foundational work on automatic differentiation, which, together with the forward gradient result in Baydin et al. [3], underpins the training substrate described in this paper. The author also thanks Don Syme, whose F# language and Units of Measure system are the type-theoretic substrate from which DTS draws its inference architecture, and whose co-authorship of the forward gradient paper connected this work to the broader AD research community. The author also thanks Paul Snively for contributions on delimited continuations and the geometric algebra problem framing documented in the companion paper [8] that informed this work.

### 13 Apr 2026 v3

@@ Abstract
Prevailing AI training infrastructure assumes reverse-mode automatic differentiation over IEEE-754 arithmetic. The memory overhead of training relative to inference, optimizer complexity, and structural degradation of geometric properties through training are consequences of this arithmetic substrate. This paper develops an alternative training architecture grounded in three prior results: the Dimensional Type System and Deterministic Memory Management framework [7], which establishes stack-eligible gradient allocation and exact quire accumulation as design-time verifiable properties; the Program Hypergraph [9], which establishes grade preservation through geometric algebra computations as a type-level invariant; and the b-posit bounded-regime design [11], which makes posit arithmetic tractable across hardware targets conventionally considered inference-only. Their composition enables depth-independent training memory bounded to approximately twice the inference footprint, grade-preserving weight updates, and exact gradient accumulation, applicable uniformly to loss-function-optimized and spike-timing-dependent neuromorphic models. We introduce Bayesian distillation, a mechanism by which the latent prior structure of a general-purpose model is extracted through the ADM training regime, resolving the data-scarcity bootstrapping problem for domain-specific training. For deployment, we introduce warm rotation, an operational pattern in which an updated model transitions into an active inference pathway without service interruption, with structural correctness formalized through PHG certificates and signed version records. The result is a class of domain-specific AI systems that are smaller and more precise than general-purpose models, continuously adaptive, verifiably correct with respect to the physical structure of their domains, and initializable from existing models.
@@ 1.2 The Three Enabling Results
The second is the Program Hypergraph established in [9]. That work demonstrated that grade in Clifford algebra is a type-level invariant under the DTS framework’s dimensional group structure, that grade inference determines the non-zero entries of the geometric product Cayley table at design time, before any arithmetic occurs, and that PHG saturation semantics correctly formalize multi-way geometric constraints that binary edge graphs cannot represent without information loss. The critical extension for training: the dual-number representation used in forward-mode autodiff augments each primal multivector with a tangent component of the same grade. PHG grade inference applies identically to tangent computations. A grade-2 bivector weight has a grade-2 gradient by the chain rule closure of grade-preserving operations, a design-time invariant discharged by the type system. A candidate update that would introduce a component at grade $j\neq k$ does not compile; the corresponding weight trajectory is structurally unreachable, and the violation is prevented at elaboration, not surfaced after the fact. In IEEE-754 the same corruption arises as silent numerical drift during accumulation and must be inferred from downstream behavior; the type system forecloses the trajectory at elaboration time and surfaces the violation in the Lattice language server as the code is written.
The third is the b-posit bounded-regime design [11]. The bounded regime field (limited to $\leq 6$ bits) reduces posit decoder hardware cost to 79% less power, 71% less area, and 60% less latency relative to standard posit decoders, achieving cost competitive with IEEE-754 float32 decoders on NPU-class hardware. This is the threshold at which posit arithmetic becomes tractable on edge and embedded hardware targets. Below this threshold, posit arithmetic required data center hardware to justify its precision advantages. At this threshold, posit arithmetic with quire accumulation is a realistic substrate for AI training on devices that also run inference.
@@ 3.4 Bayesian Distillation from General-Purpose Models
Recent empirical work [25] establishes the key precondition: latent Bayesian structure in general-purpose LLMs is accessible through targeted fine-tuning and generalizes across domains. This result, developed independently of the ADM framework, confirms that the prior knowledge absorbed during large-scale pretraining is not merely a surface statistical regularity but a structural property that can be elicited and directed. The van Steenkiste and Linzen study evaluates belief updating on a sequential recommendation task with a five-round horizon and discrete preference structure, and their generalization claim covers transfer across semantically related task families of comparable structure. Physical domains of the kind targeted by ADM, with continuous dimensional parameters, geometric grade structure, and strong multi-physics coupling, lie outside the evaluated task family. The result therefore establishes that elicitation is empirically feasible in at least one non-trivial regime; whether the same elicitation procedure transfers cleanly to dimensionally annotated physical domains is an open question, and the filtering stage described below is designed in part to absorb the transfer gap.
@@
From type constraints to weight-space priors.
The filtering stage is given a concrete construction through the mapping between type-level constraints on a layer’s interface and the support of the prior on that layer’s weight matrix. Consider a linear layer $f(x)=Wx$ with input dimensional type $T_{\mathrm{in}}$ (say, velocity, $\mathrm{m\,s^{-1}}$) and output dimensional type $T_{\mathrm{out}}$ (say, pressure gradient, $\mathrm{Pa\,m^{-1}}$). DTS inference [7] assigns $W$ the dimensional signature $T_{\mathrm{out}}/T_{\mathrm{in}}=\mathrm{Pa\,s\,m^{-2}}$. Any weight configuration with entries that do not carry this dimension is outside the admissible subspace $\mathcal{W}_{\mathrm{dim}}\subset\mathbb{R}^{d_{\mathrm{out}}d_{\mathrm{in}}}$ (the latter denoting the $d_{\mathrm{out}}\times d_{\mathrm{in}}$ real matrix space). For a Clifford layer with input grade $k$ and output grade $k^{\prime}$, PHG inference [9] restricts $W$ to the subspace $\mathcal{W}_{\mathrm{grade}}$ of linear maps between the grade-$k$ and grade-$k^{\prime}$ subspaces of the ambient multivector space; entries at other grade pairs are type violations and carry zero probability. The admissible weight support is the intersection $\mathcal{W}_{\mathrm{adm}}=\mathcal{W}_{\mathrm{dim}}\cap\mathcal{W}_{\mathrm{grade}}$.
The type system therefore induces a prior of the form
$p(W)\propto\mathbf{1}[W\in\mathcal{W}_{\mathrm{adm}}]\cdot\tilde{p}(W),$
where $\tilde{p}(W)$ is a density over $\mathcal{W}_{\mathrm{adm}}$ and the indicator is exact, not a penalty. Writing $\Pi_{\mathrm{adm}}$ for the orthogonal projector onto $\mathcal{W}_{\mathrm{adm}}$, the filtering stage maps a distillation candidate $W_{\mathrm{cand}}$ drawn from the LLM’s extracted distribution to $\Pi_{\mathrm{adm}}W_{\mathrm{cand}}$, and reweights the density of the surviving mass to match $\tilde{p}$. The density $\tilde{p}$ is constructed from the distilled distribution by conditioning on $\mathcal{W}_{\mathrm{adm}}$: the relative probabilities the LLM assigned to compatible weight configurations are retained, while mass on incompatible configurations is removed rather than penalized. This is the operational sense in which the type constraints act as a filter rather than a regularizer: the regularizer formulation permits incompatible configurations at high energy cost, while the filter formulation assigns them zero probability and removes them from the support of the prior.
Two consequences follow. First, the bridge is mechanical, not heuristic: the admissible support is determined at elaboration time by type inference, and the projector $\Pi_{\mathrm{adm}}$ is a deterministic function of the layer’s type signature. Second, the distilled prior is a proper probability distribution on $\mathcal{W}_{\mathrm{adm}}$ whose posterior under domain data is well-defined and computable by the forward-mode procedure of Section 3.1. The distillation regime inherits the calibration properties of the LLM where those properties are consistent with the type system, and discards them where they are not.
@@ 4.2 Memory Feasibility Across Hardware Targets
The $2\times$ bound covers point-estimate (MAP) training. The full posterior approximation of Section 3.1 maintains a low-rank estimate of the Hessian’s leading eigenvectors in addition to the primal and tangent slots, for a total memory cost of $(2+r/d)\times$ the inference footprint, where $d$ is the parameter dimension and $r$ is the retained rank. For the rank schedule considered here ($r\ll d$, typically $r$ in the tens to low hundreds and independent of depth), the additional term is small relative to the primal and tangent components, and the qualitative bound remains constant-multiple of inference memory. Reverse-mode Bayesian approximations with equivalent curvature information require this rank on top of the activation tape, so the asymptotic separation in model depth is preserved.
@@ 5.1 Grade Preservation Through Training
Proof sketch. The dual-number extension of a grade-$k$ primal is a grade-$k$ tangent. PHG grade inference applies identically to the tangent computation as to the primal. The forward-mode procedure of [3] samples a random projection direction $v$ per step; under the PHG type system, $v$ is drawn from the grade-$k$ subspace of the weight space, not from the full multivector space. This constraint is discharged at elaboration time as a consequence of the grade type of $W$: the tangent slot associated with $W$ has the same grade as $W$ itself, and the projection vector that populates it inherits that grade by type. The directional derivative $\langle\nabla_{W}\mathcal{L},v\rangle$ is then a scalar (grade-0) result, computed as the quire accumulation of products between the grade-$k$ weight tangent components and the grade-$k$ projection $v$. The gradient estimate $\nabla_{W}\mathcal{L}$ is the reconstruction of the weight-space vector from the scalar projection estimates, and lies in the grade-$k$ subspace by construction because the estimator is a linear combination of grade-$k$ projection vectors. Any component at grade $j\neq k$ is a type violation caught at design time. $\square$
@@ 10. Conclusion
Loss-function gradient descent over Clifford algebra neural networks and spike-timing-dependent plasticity over neuromorphic spiking networks are both instances of this architecture, and the warm rotation mechanism and versioning infrastructure of Section 4 applies uniformly to each. Trust in the resulting system is a structural property discharged at design time by the elaboration of the program’s type and grade invariants and anchored in cryptographically signed version records, and is established without reliance on benchmark performance against a curated test set.
@@ Acknowledgments
The author thanks John L. Gustafson for guidance on posit arithmetic, the b-posit bounded-regime arithmetic pathway, and for Every Bit Counts: Posit Computing [10], whose Chapter 13 provides the definitive parameterization reference for application-specific posit systems. The treatment of representation selection for training arithmetic in Sections 3 and 4 reflects his technical influence directly. The author thanks Barak A. Pearlmutter for his encouragement of this research program and for his foundational work on automatic differentiation, which, together with the forward gradient result in Baydin et al. [3], underpins the training substrate described in this paper. The author also thanks Don Syme, whose F# language and Units of Measure system are the type-theoretic substrate from which DTS draws its inference architecture, and whose co-authorship of the forward gradient paper connected this work to the broader AD research community. The author also thanks Paul Snively for contributions on delimited continuations and the geometric algebra problem framing documented in the companion paper [9] that informed this work. The author also thanks Martín Coll, whose Inet MLIR dialect [4] is the rewrite-rule substrate on which the program hypergraph construction of the companion paper [9] is primarily informed, and whose subsequent work on the hypergraph interchange format [5] provides the structural serialization precedent for the PHG’s cross-tool provenance handling.